OpenStudy (anonymous):

One week Joe worked 38 1/2 hours and earned $115 in tips. If his weekly earnings totaled $275.85, what is his hourly rate?

pooja195 (pooja195):

plz take a look

OpenStudy (anonymous):

thanksz :)

OpenStudy (anonymous):

ok first look at what the weekly pay was and take away the tips (2.75.85-115=60.85). Then divide that by the hours worked that week (60.85/38.5hrs). This gives you the hourly rate for the week of 1.580519481 or rounded to the nearest penny it is $1.58.

OpenStudy (anonymous):

I think I did that based on 175.85. It should have been 275.85 but you could still do everything else the same way and get the answer. Just go to the like they provided.

OpenStudy (anonymous):

275.85-115=160

OpenStudy (anonymous):

then 160/38.5hrs=about $4.16 (this is rounded to the nearest penny from 4.155844156)

OpenStudy (anonymous):

can u help me w this one, Bookkeeper - Annual salary: $29,400; What is his monthly salary?

OpenStudy (anonymous):

29400/12=2450

OpenStudy (anonymous):

12 comes from 12 months in the year

pooja195 (pooja195):

Bookkeeper monthly salary would be $2450.

pooja195 (pooja195):

u could do this on a calculatro

OpenStudy (anonymous):

oo okaii :)

pooja195 (pooja195):

salary would be $2450

OpenStudy (anonymous):

w abt this one. Swimming coach - Weekly pay: $143.75; What is her annual salary?

OpenStudy (anonymous):

Use the formula above.weekly pay/12

OpenStudy (anonymous):

scratch that. it's weekly pay times however many week are in a year (52, hint hint)